# Elephants Never Forget: Memorization and Learning of Tabular Data in Large Language Models

### v1

@@ 4.4 Language models have a limited ability to act as statistical predictors
In the previous section, we have seen that there is a significant performance drop between task and statistical, indicating that the few-shot learning performance of LLMs significantly relies on the LLM’s world knowledge (Yu et al., 2023). In this section, we study the ability of language models to act as in-context statistical learners. This relates to works that have used small, specifically trained language models to show that in-context learning can learn simple function classes, akin to supervised learning (Garg et al., 2022). We now investigate this question in the context of GPT-3.5 and GPT-4.
@@ 5 Drawing Random Samples from Datasets seen during Training
Figure 4 depicts the longitude and latitude of random samples on the California Housing dataset. To draw these samples, we conditioned GPT-3.5 with samples from other datasets, revealing no information about the feature values on the California Housing dataset. Interestingly, the samples drawn by the LLM follow the overall summary statistics in the data. For example, the feature correlations of the samples match the feature correlation in the original dataset (Supplement Figure 9). At the same time, the samples are not copied from the training data (Supplement Table 8). In this sense, it can be said that GPT-3.5 generates novel random samples from the dataset.
@@
Figure 4: GPT-3.5 can draw random sample from the California Housing dataset. We only provide the model with the name of the dataset and the feature names. The diversity of the generated samples depends on the temperature parameter. For small temperatures, the samples are concentrated around the mode of data. As temperature increases, the samples become more diverse and similar to the data distribution. At large temperatures, some samples lie outside the support of the data distribution. The reader might want to compare with Figure 1 in (Borisov et al., 2023).
@@ 6 Related Work
Recent works have demonstrated the capabilities of LLMs on tasks with tabular data (Dinh et al., 2022; Narayan et al., 2022; Vos et al., 2022; Wang et al., 2023; McMaster et al., 2023). In particular, Borisov et al. (2023) and Hegselmann et al. (2023) have shown that LLMs can be fine-tuned to generate and classify tabular data. The phenomenon of in-context learning has attracted much research interest (Brown et al., 2020; Bhattamishra et al., 2023; Wei et al., 2023; Li et al., 2023). It has been shown that the forward pass of a transformer can mimic gradient descent (Von Oswald et al., 2023) and that LLMs can learn simple function classes in-context (Garg et al., 2022; Bhattamishra et al., 2023).
Data contamination is studied in many LLM reports and research papers, often adopting an n-gram-based definition (OpenAI, 2023; Nori et al., 2023; Touvron et al., 2023). Jiang et al. (2024) investigate data contamination in GPT-2 by adding benchmark datasets to the pre-training data. Yang et al. (2023) study the effects of rephrasing widely used benchmarks.
Membership inference attacks are a set of methods to detect if a model has seen a text during training (Mahloujifar et al., 2021; Choquette-Choo et al., 2021; Carlini et al., 2022a; Mireshghallah et al., 2022; Shi et al., 2024; Mattern et al., 2023). Recently, Duan et al. (2024) question the value of membership inference attacks for large language models. Shi et al. (2024) introduce the MIN-K%-PROB method to detect pre-training data from LLMs.
@@
LLMs tend to perform better on tasks more likely to occur in the training data (Wu et al., 2023; McCoy et al., 2023; Dziri et al., 2023b). Our work complements this emerging literature by (1) rigorously demonstrating that GPT-3.5 and GPT-4 have seen certain tasks frequently during training (whereas previous works have relied on heuristics) and (2) showing that LLMs overfit on uncommon tasks that were seen frequently during training. [^5] We would not consider “the chemical properties of wines grown in the same region in Italy but derived from three different cultivars” – the task in the Wine dataset – a common task.
@@ 9 Reproducibility Statment
The dataset transforms, as well as the memorization tests, can be performed with the open source package tabmemcheck.
@@ Appendix D Memorization Tests
This section gives additional details on the different tests and the quantitative results of the memorization tests. As a general observation, one might expect the results of the tests to be somewhat ambiguous, especially the quantitative memorization tests. However, we find that the results of the different tests were surprisingly clear-cut, with only very few ambiguous cases.
The main prompt strategy of the different tests is to condition the model with few-shot examples from other datasets. For example, we provide the model with a chat history where it provides the feature names on other datasets and then ask it to complete the feature names on the test dataset. We use the Iris, Adult Income, Kaggle Titanic, Wine, and California Housing datasets for the few-shot examples (when we are testing a dataset, it is, of course, not part of the few-shot examples).
@@
D.1 Feaure Names
The feature names test uses our general prompt structure with few-shot examples, given in Figure 7. In all cases, the result of this test was unambiguous. The model would either list the names of the different features in the same format as in the CSV file of the dataset, or it would offer clearly made-up feature names.
D.2 Feature Values
To test whether the model is able to model the feature values of the data, we ask it to provide samples from the dataset. We then check whether the formatting of the sampled feature values is the same as the formatting in the training data.
D.3 Feature Distribution
D.4 Conditional Distribution
D.5 Header Test
We split the dataset at random positions in rows 2, 4, 6, and 8 and ask the model to complete the dataset from the given position. We condition the model on this task with few-shot examples and consider the best completion. The test succeeds if the model completes at least the next row in the dataset. This was unambiguous on all datasets (the model completed many rows or not even a single row). The prompt is depicted in Figure 6.
D.6 Row Completion, Feature completion, and First Token Test
The prompts for the row completion and first token test are similar to the header test. Table 4 depicts the quantitative results of these tests. Table 5 depicts the detailed results of the first token test.
@@ D.6 Row Completion, Feature completion, and First Token Test
Table 4: Detailed results of the row completion and feature completion tests from Table 2 in the main paper. The table depicts the number of rows and features that were correctly completed.
@@
Table 5: Detailed results of the first token test from Table 2 in the main paper. The table depicts the number of first tokens that were correctly completed and the overall accuracy of the completions (in brackets). The last row depicts the completion accuracy that can be reached using traditional statistical predictors (this improves upon the mode if the rows are not random).
@@ Appendix E Dataset Formats
Perturbed. Starting from the original data, we perturb continuous and discrete variables by randomly perturbing individual digits. We perturb the data so that we (1) change at least a single digit in all non-zero values in the data and (2) perturb with a relative error of approximately 1%. This means that the perturbation depends on a variable’s magnitude and number of significant digits. The perturbations do not affect the format of the variables in the data, meaning that the perturbed data still “look like” the original data. However, the numerical values are slightly different.
@@
Appendix F Prompts

### v3

@@ 5 Drawing Random Samples from Datasets seen during Training
Figure 4: GPT-3.5 can draw random samples from the California Housing dataset. We only provide the model with the name of the dataset and the feature names. The diversity of the generated samples depends on the temperature parameter. For small temperatures, the samples are concentrated around the mode of data. As temperature increases, the samples become more diverse and similar to the data distribution. At large temperatures, some samples lie outside the support of the data distribution. The reader might want to compare with Figure 1 in Borisov et al. (2023).
@@
Figure 4(d) depicts the longitude and latitude of random samples on the California Housing dataset. To draw these samples, we conditioned GPT-3.5 with samples from other datasets, revealing no information about the feature values on the California Housing dataset. Interestingly, the samples drawn by the LLM follow the overall summary statistics in the data. For example, the feature correlations of the samples match the feature correlation in the original dataset (Supplement Figure 16). At the same time, the samples are not copied from the training data (Supplement Table 11). In this sense, it can be said that GPT-3.5 generates novel random samples from the dataset.
@@ 6 Related Work
Recent works have demonstrated the capabilities of LLMs on tasks with tabular data (Dinh et al., 2022; Narayan et al., 2022; Vos et al., 2022; Wang et al., 2023; McMaster et al., 2023). In particular, Borisov et al. (2023) and Hegselmann et al. (2023) have shown that LLMs can be fine-tuned to generate and classify tabular data. The phenomenon of in-context learning has attracted much research interest (Brown et al., 2020; Wei et al., 2023; Li et al., 2023). It has been shown that the forward pass of a transformer can mimic gradient descent (Von Oswald et al., 2023) and that LLMs can learn simple function classes in-context (Garg et al., 2022; Bhattamishra et al., 2024).
Data or task contamination is studied in many LLM reports and research papers, often adopting an n-gram-based definition (OpenAI, 2023; Nori et al., 2023; Touvron et al., 2023). Jiang et al. (2024) investigate data contamination in GPT-2 by adding benchmark datasets to the pre-training data. Yang et al. (2023) study the effects of rephrasing widely used benchmarks. Li & Flanigan (2024) analyze how the few-shot learning performance of LLMs varies between datasets released before and after the training data was collected.
Membership inference attacks are a set of methods to detect if a model has seen a text during training (Mahloujifar et al., 2021; Choquette-Choo et al., 2021; Carlini et al., 2022a; Mireshghallah et al., 2022; Shi et al., 2024; Mattern et al., 2023). Recently, Duan et al. (2024) question the value of membership inference attacks for large language models. Shi et al. (2024) introduce the Min-KProb method to detect pre-training data from LLMs.
@@
LLMs tend to perform better on tasks more likely to occur in the training data (Wu et al., 2023; McCoy et al., 2023; Dziri et al., 2023b). Our work complements this emerging literature by (1) rigorously demonstrating that GPT-3.5 and GPT-4 have seen certain tasks frequently during training (whereas previous works have relied on heuristics) and (2) showing that LLMs overfit on uncommon tasks that were seen frequently during training.
@@ 9 Reproducibility Statment
Code to replicate the results in this paper is available at https://github.com/interpretml/LLM-Tabular-Memorization-Checker. This includes, in particular, the different memorization tests and the dataset transformations.
@@ Appendix D Memorization Tests
This Section provides additional details on the memorization tests and their quantitative results. We also detail a number of measures that assess what the LLM knows about the dataset (for example, the feature names).
With regard to the interpretation of the memorization tests, it is important to note that the quantitative test results (e.g., how many rows the LLM managed to complete in the row completion test) are not sufficient to judge memorization. Instead, one has to consider the completion rate of the LLM and the amount of entropy in the dataset. For example, the OpenML Diabetes dataset contains an individual’s glucose level, blood pressure, and BMI, as well as other measurements that are at least in part random and highly unique. If an LLM can consistently generate even a few rows of this dataset, it is strong evidence of memorization. To give another example, the Iris dataset contains many rows that are near-duplicates. This means that an LLM might also achieve a non-zero row completion rate by chance or prediction. Because of this complication, we manually judge the results in Table 2 and Table 3 in the main paper, considering the quantitative results of the different tests and the structure of the dataset.
As a general observation, one might expect the results of the tests to be somewhat ambiguous. It turns out, however, that GPT-3.5 and GPT-4 (as well as other LLMs) have the remarkable property of regurgitating quite long and complex strings from their training data. For example, GPT-4 is able to generate the entire Iris dataset, consisting of 151 rows, given only the first couple of rows as prompt. Thus, while there are few cases of ambiguity (marked as question marks in the respective tables), the evidence of memorization is usually fairly clear (and conversely, there are cases where the model gets the dataset’s content completely wrong).
@@
D.1 Qualitative Examples
Figures 5-10 on the following pages illustrate the different memorization tests. They depict the Leveshtein string distance between the model response and the actual content in the dataset. For the different tests, we depict two cases: one with evidence of memorization and one without.
Figure 5: Header Test on Adult Income. The LLM is prompted with the first couple of rows of the dataset (black text) and responds with the next 7 rows of the dataset, exactly as they appear in the CSV file of the dataset (green text). The text color illustrates the Levenshtein string distance between the text in the CSV file and the model response. An entirely green row means that the model responded with the next row exactly as it occurs in the CSV file. Red color indicates that the model made a mistake, and violet color indicates that the model missed a digit. In this particular example, the model response is all green because it is equal to the content of the CSV file. Best viewed in digital format.
Figure 6: Row Completion Test on Iris. The LLM is prompted with the previous rows of the dataset and responds with the next row. The figure depicts the Levenshtein string distance between the model responses and the actual next rows in the CSV file. An entirely green row means that the model responded with the next row exactly as it occurs in the CSV file. Red color indicates that the model made a mistake, and violet color indicates that the model missed a digit. Best viewed in digital format.
Figure 7: Header Test on ACS Income. The LLM is prompted with the first couple of rows of the dataset (black text) and responds with 8 more rows. The text color illustrates the Levenshtein string distance between the text in the CSV file and the model response. An entirely green row means that the model responded with the next row exactly as it occurs in the CSV file. Red color indicates that the model made a mistake, and violet color indicates that the model missed a digit. Best viewed in digital format.
Figure 8: Row Completion Test on FICO. The LLM is prompted with the previous rows of the dataset and responds with the next row. The text color illustrates the Levenshtein string distance between the text in the CSV file and the model response. An entirely green row means that the model responded with the next row exactly as it occurs in the CSV file. Red color indicates that the model made a mistake, and violet color indicates that the model missed a digit. Best viewed in digital format.
Figure 9: Feature Completion Test on OpenML Diabetes. The LLM is prompted with the feature values of a row, except for a single missing feature. The model responds with the value of the missing feature. The figure depicts the Levenshtein string distance between the model responses and the actual feature values in the CSV file. An entirely green text means that the model responded with the feature value exactly as it occurs in the CSV file. Red color indicates that the model made a mistake, and violet color indicates that the model missed a digit. Best viewed in digital format.
Figure 10: Feature Completion Test on Adult Income. The LLM is prompted with the feature values of a row, except for a single missing feature. The model responds with the value of the missing feature. The figure depicts the Levenshtein string distance between the model responses and the actual feature values in the CSV file. An entirely green text means that the model responded with the feature value exactly as it occurs in the CSV file. Red color indicates that the model made a mistake, and violet color indicates that the model missed a digit. Best viewed in digital format.
@@ D.1 Qualitative Examples
Table 5: Quantitative test results of the row completion and feature completion tests from Table 2 in the main paper. The table depicts the number of rows and features that were correctly completed.
@@
Table 6: Quantitative test results of first token test from Table 2 in the main paper. The table depicts the number of first tokens that were correctly completed and the overall accuracy of the completions (in brackets). The last row depicts the completion accuracy that can be reached using traditional statistical predictors (this improves upon the mode if the rows are not random).
D.2 Quantitative Test Results
Table 5 and Table 6 depict the quantitative results of the row completion, feature completion, and first token test.
For the row completion test depicted in Table 5, we count the number of times that a row is correctly completed. In comparison with Figure 6 and Figure 8, which depict model responses on the row completion task, this means that we count the number of entirely green rows (i.e. zero Levenshtein distance between the row in the CSV file and the model response). From the numbers depicted in Table 5, we see that the number of row completions is either zero / in very low digits or fairly large, with over 90% correctly completed rows on some datasets. As mentioned before, the number of correct row completions has to be judged in comparison with the amount of entropy in the data. A concrete example of this is depicted in Figure 8, where the correctly completed row on FICO is one of the many duplicate rows in the dataset that consists repeatedly of the number -9. Of course, this does not provide evidence of memorization.
For the feature completion test in Table 5, we count the number of times that a feature value is correctly completed. We chose the feature to be completed as a highly unique feature (the Iris dataset does not contain such a feature, so we don’t conduct the test on this dataset). The results obtained for the feature completion test are very similar to those of the row completion test. Again, we see that the number of feature completions is either zero / in very low digits or fairly large, with over 90% correctly completed feature values on some datasets.
For the first token test depicted in Table 6, we count the number of times that the model correctly predicts the first token of the next row in the CSV file. The prompts for this test are the same as for the row completion test; the difference is that we don’t assess whether the entire row is correctly completed but only consider the first token. The first token test is especially interesting for the Adult dataset because we know from the construction of the dataset that the rows are ordered at random. In this case, a model that has never seen the dataset cannot be better at completing the first token than the mode. Interestingly, we see that the completion rate of GPT-4 is exactly at the rate of the mode.
D.3 Additional Test Descriptions
Here we provide additional details on the different tests.
D.3.1 Feature Names
The feature names test asks the LLM to complete the names of the features of a dataset, given the name of the first feature. The test uses our general prompt structure with few-shot examples, given in Figure 14. In all cases, the result of this test was unambiguous. The model would either list the names of the different features in the same format as in the CSV file of the dataset, or it would offer clearly made-up feature names.
D.3.2 Feature Values
The feature values test asks the LLM to provide observations from the dataset that have valid feature values (most importantly, valid string values for categorical features). This test uses the prompt structure used in Section 5 in the main paper to draw samples from the dataset. We then manually check whether the formatting of the sampled feature values is the same as the formatting in the training data.
D.3.3 Feature Distribution
D.3.4 Conditional Distribution
D.3.5 Header Test
For the header test, we split the dataset at random positions in rows 2, 4, 6, and 8 and ask the model to complete the dataset from the given position. We condition the model on this task with few-shot examples and consider the best completion. The test succeeds if the model completes at least the next row in the dataset. This was unambiguous on all datasets (the model completed many rows or not even a single row). The prompt is depicted in Figure 13.
@@ Appendix E Dataset Formats
Perturbed. Starting from the original data, we perturb continuous and discrete variables by randomly perturbing individual digits. We perturb the data so that we (1) change at least a single digit in all non-zero values in the data and (2) perturb with a relative error of approximately 1%. This means that the perturbation depends on a variable’s magnitude and number of significant digits. The perturbations do not affect the format of the variables in the data, meaning that the perturbed data still “looks like” the original data. However, the numerical values are slightly different.
@@
Appendix F Replication with Llama 3.1 70B and Gemma 2 27B
Panel A Titanic Adult Diabetes Wine Iris
Original $0.81_{\scriptscriptstyle.01}$ $0.77_{\scriptscriptstyle.01}$ $0.69_{\scriptscriptstyle.02}$ $0.92_{\scriptscriptstyle.02}$ $0.97_{\scriptscriptstyle.01}$
Perturbed $0.79_{\scriptscriptstyle.01}$ $0.78_{\scriptscriptstyle.01}$ $0.69_{\scriptscriptstyle.02}$ $0.91_{\scriptscriptstyle.02}$ $0.93_{\scriptscriptstyle.02}$
Task $0.78_{\scriptscriptstyle.01}$ $0.72_{\scriptscriptstyle.01}$ $0.70_{\scriptscriptstyle.02}$ $0.87_{\scriptscriptstyle.03}$ $0.94_{\scriptscriptstyle.02}$
Statistical $0.63_{\scriptscriptstyle.02}$ $0.72_{\scriptscriptstyle.01}$ $0.66_{\scriptscriptstyle.02}$ $0.85_{\scriptscriptstyle.03}$ $0.91_{\scriptscriptstyle.02}$
LR / GBT 0.79   /   0.80 0.86   /   0.87 0.78   /   0.75 0.98   /   0.96 0.97   /   0.95
Panel B S. Titanic ACS Income ICU FICO ACS Travel
Original $0.69_{\scriptscriptstyle.01}$ $0.78_{\scriptscriptstyle.01}$ $0.69_{\scriptscriptstyle.05}$ $0.69_{\scriptscriptstyle.01}$ $0.62_{\scriptscriptstyle.02}$
Perturbed $0.69_{\scriptscriptstyle.01}$ $0.78_{\scriptscriptstyle.01}$ $0.71_{\scriptscriptstyle.05}$ $0.68_{\scriptscriptstyle.01}$ $0.62_{\scriptscriptstyle.02}$
Task $0.71_{\scriptscriptstyle.01}$ $0.77_{\scriptscriptstyle.01}$ $0.71_{\scriptscriptstyle.05}$ $0.63_{\scriptscriptstyle.02}$ $0.61_{\scriptscriptstyle.02}$
Statistical $0.66_{\scriptscriptstyle.01}$ $0.59_{\scriptscriptstyle.02}$ $0.57_{\scriptscriptstyle.05}$ $0.61_{\scriptscriptstyle.02}$ $0.49_{\scriptscriptstyle.02}$
LR / GBT 0.78   /   0.78 0.80   /   0.80 0.76   /   0.66 0.70   /   0.69 0.64   /   0.67
Table 7: Few-shot learning performance of LLama 3.1 70B across different tabular datasets. This table replicates Table 4 in the main paper.
Panel A Titanic Adult Diabetes Wine Iris
Original $0.81_{\scriptscriptstyle.01}$ $0.82_{\scriptscriptstyle.01}$ $0.70_{\scriptscriptstyle.02}$ $0.93_{\scriptscriptstyle.02}$ $0.99_{\scriptscriptstyle.01}$
Perturbed $0.80_{\scriptscriptstyle.01}$ $0.82_{\scriptscriptstyle.01}$ $0.71_{\scriptscriptstyle.02}$ $0.92_{\scriptscriptstyle.02}$ $0.92_{\scriptscriptstyle.02}$
Task $0.79_{\scriptscriptstyle.01}$ $0.78_{\scriptscriptstyle.01}$ $0.72_{\scriptscriptstyle.02}$ $0.90_{\scriptscriptstyle.02}$ $0.92_{\scriptscriptstyle.02}$
Statistical $0.64_{\scriptscriptstyle.02}$ $0.73_{\scriptscriptstyle.01}$ $0.67_{\scriptscriptstyle.02}$ $0.88_{\scriptscriptstyle.02}$ $0.88_{\scriptscriptstyle.03}$
LR / GBT 0.79   /   0.80 0.86   /   0.87 0.78   /   0.75 0.98   /   0.96 0.97   /   0.95
Panel B S. Titanic ACS Income ICU FICO ACS Travel
Original $0.68_{\scriptscriptstyle.01}$ $0.77_{\scriptscriptstyle.01}$ $0.73_{\scriptscriptstyle.04}$ $0.64_{\scriptscriptstyle.02}$ $0.63_{\scriptscriptstyle.02}$
Perturbed $0.68_{\scriptscriptstyle.01}$ $0.77_{\scriptscriptstyle.01}$ $0.70_{\scriptscriptstyle.05}$ $0.64_{\scriptscriptstyle.02}$ $0.64_{\scriptscriptstyle.02}$
Task $0.64_{\scriptscriptstyle.02}$ $0.76_{\scriptscriptstyle.01}$ $0.75_{\scriptscriptstyle.04}$ $0.63_{\scriptscriptstyle.02}$ $0.59_{\scriptscriptstyle.02}$
Statistical $0.65_{\scriptscriptstyle.02}$ $0.59_{\scriptscriptstyle.02}$ $0.54_{\scriptscriptstyle.05}$ $0.54_{\scriptscriptstyle.02}$ $0.54_{\scriptscriptstyle.02}$
LR / GBT 0.78   /   0.78 0.80   /   0.80 0.76   /   0.66 0.70   /   0.69 0.64   /   0.67
Table 8: Few-shot learning performance of Gemma 2 across different tabular datasets. This table replicates Table 4 in the main paper.
In this Section, we replicate the experiments from Section 4 in the main paper using Llama 3.1 70B and Gemma 2 27B (Dubey et al., 2024; Riviere et al., 2024). We do not perform any changes to the setup of the experiment, that is we send the same prompts to the open-weight models that were previously send to GPT-3.5 and GPT-4 (including the ordering of the few-shot examples). We use https://www.together.ai/ as our inference engine, which means that the experiments can be performed with the same code that was used to send the prompts to the OpenAI API. The model endpoints are Meta-Llama-3.1-70B-Instruct-Turbo and google/gemma-2-27b-it. Code is available at https://github.com/interpretml/LLM-Tabular-Memorization-Checker.
While the experiment in this Section can be seen as a replication of the experiment in Section 4 in the main paper, there is an important difference between the experiment detailed in this Section and the experiment in the main paper. Namely, the datasets that are ”novel” for the versions of GPT-3.5 and GPT-4 considered in Section 4 might have been included in the training data of the open-weight models. Moreover, as discussed in Section 3 in the main paper, the patterns of dataset memorization are different for the open-weight LLMs: Unlike GPT-3.5 and GPT-4, the open-weight LLMs do not seem to have memorized entire datasets, with the exception of the Iris dataset. Naturally, these two points are very important for the interpretation of the results, since we might not necessarily expect the same results if a dataset was included in the pre-training data, or if it was not memorized by the model.
With these caveats in mind, let us consider the few-shot learning results for Llama 3.1 70B and Gemma 2 27B. The results for Llama 3.1 are depicted in Table 7, and the results for Gemma 2 are depicted in Table 8. Figure 11 summarizes the results in both tables.
Interestingly, the results for Llama 3.1 70B are very similar to the results obtained with GPT-3.5 (compare Figure 11 and Figure 2 in the main paper). This is especially true on the popular tabular datasets, where both models show a similar decaying pattern across dataset transformations and also similar absolute performance (this is depicted in Subfigure (a) of Figure 11). On the datasets that are novel for GPT-3.5, both models show a similar pattern across dataset transformations, but Llama 3.1 70B performs, on average, better.
The results for Gemma 2 across the popular tabular datasets are similar to the results in the main paper, that is we see the characteristic decaying pattern across dataset transformations (this is depicted in Subfigure (c) of Figure 11). Gemma 2 depicts a slightly decaying pattern across dataset transformations also for the datasets that are novel for GPT-3.5 and GPT-4, even if this pattern is less pronounced and not statistically significant (this is depicted in Subfigure (d) of Figure 11). This result might be due to random fluctuations, or it might be that some of the datasets are included in the pre-training data of the model. More research would be needed to understand this in detail.
At a high level, it is interesting to see that the results obtained with Llama 3.1 70 B and Gemma 2 27B are overall similar to the results obtained with GPT-3.5 and GPT-4, even if some of the assumptions around the training data cutoff and memorization don’t necessarily hold in the same way. Of course, this might be because the popular tabular datasets are still more likely to be included in the pre-training data than some of the novel datasets, even if the novel datasets were, strictly speaking, released before the cutoff of the training data.
(a) Llama Panel A
(b) Llama Panel B
(c) Gemma Panel A
(d) Gemma Panel B
Figure 11: Few-shot learning performance of Llama 3.1 70B and Gemma 2 27B across the datasets depicted in Panel A and Panel B of Table 7 and Table 8, respectively. This figure replicates Figure 2 in the main paper. Note that we have replaced ”memorized” and ”novel” in the description of the Figure with ”Panel A” and ”Panel B”. This is because the results of the memorization tests for Llama 3.1 70B and Gemma 2 27B indicate that there is less memorization than for GPT-3.5 and GPT-4.
Appendix G Prompts